### v1

# Plasmonic metamaterial time crystal

Abstract
Periodically driven optical materials and metamaterials have recently emerged as a promising platform for realizing photonic time crystals (PTCs)—systems whose optical properties are strongly and periodically modulated on time scales comparable to the optical cycle of light Asgari_2024 . These time-varying structures are the temporal counterparts of spatial photonic crystals (SPCs), for which a large and periodic dielectric contrast is achieved spatially on wavelength scales. Just as SPCs have revolutionized control over light–matter interactions by engineering the photonic density of states in space Yablonovitch_1987 ; John_1987 ; Joannopoulos_2011 ; Lu_2014 , PTCs promise comparable breakthroughs from a fundamentally new perspective: a temporal one Lustig_2018 ; Lyubarov_2022 ; Dikopoltsev_2022 ; Li_2023 ; Wang_2024 ; Park_2025 . However, harnessing such phenomena at optical frequencies poses severe experimental challenges Asgari_2024 ; Hayran_2022 ; Saha_2023 ; Hayran_2025 , as it requires order-unity modulation depths of the optical properties at optical cycle rates, a regime that has remained elusive to date.
Here, we report the first optical realization of a photonic time crystal, achieved with a surface plasmon cavity metamaterial operating at Terahertz frequencies. We demonstrate strong (near-unity) and coherent (sub-optical cycle) periodic driving of the plasmonic metamaterial enabled by field-induced dynamical modulation of the carriers’ kinetic energy and effective mass—reaching up to 80% of their rest mass, an exceptionally high value that forms the basis for time crystalline phenomena with plasmons Kiselev_2024 ; Feinberg_2025 . Our experimentally informed theory reveals rich physics within the experimentally accessible parameter regime of this system, including parametric amplification and entangled plasmon generation, and establishes a robust new platform for time-domain photonics.
Introduction
[…]
Systems subjected to periodic temporal modulations, known as Floquet systems, have been extensively studied across diverse platforms, including mechanical Trainiti_2016 , atomic Eckardt_2017 , and solid-state Wang_2013 ; Nova_2016 ; Cavalleri_2017 ; McIver_2019 ; Oka_2019 ; Rudner_2020 . In such systems, periodic driving is known to provide fascinating properties not accessible in equilibrium Bukov_2015 and the ability to tailor emergent properties through Floquet engineering of effective Hamiltonians Goldman_2014 ; Weitenberg_2021 . Over the past few years, these ideas have increasingly influenced the area of photonics. Periodic temporal modulation of optical systems enables time, in addition to space, to be utilized as a synthetic dimension for the tailoring of optical eigenmodes and response functions, an approach described within the framework of four-dimensional or Floquet metamaterials Engheta_2020 ; Yin_2022 ; Galiffi_2022 ; Sapienza_2023 ; Engheta_2023 ; Horsley_2023 ; Sisler_2024 . For spatially homogeneous yet ultrafast temporal modulations approaching the timescale of the optical cycle, the related concept of photonic time crystals (PTCs) has recently emerged Asgari_2024 ; Lustig_2018 ; Lyubarov_2022 ; Dikopoltsev_2022 ; Li_2023 ; Wang_2024 ; Park_2025 ; Kiselev_2024 ; Feinberg_2025 ; Pan_2023 ; Hayran_2024 ; Khurgin_2024 . As depicted in Figure 1a, PTCs can be viewed as the temporal analogues of spatial photonic crystals (SPCs) Yablonovitch_1987 ; John_1987 ; Joannopoulos_2011 ; Lu_2014 by permuting space and time coordinates on one hand and momentum and frequency on the other. In SPCs, spatial periodicity generates band replicas along the momentum axis and opens frequency gaps ($\omega$-gaps), whereas in PTCs, temporal periodicity produces replicas along the frequency axis and opens momentum gaps ($k$-gaps). Wave-behavior within or in the vicinity of these gaps is key in governing light–matter interactions and differ fundamentally between the two cases. Inside the gaps, wave solutions reduce to exponentially decaying or growing modes — spatially in SPCs and temporally in PTCs (Fig. 1a). In SPCs, only the decaying solution is physically allowed, a consequence of their time-invariant and energy-conserving nature, leading to inhibition of spontaneous emission Yablonovitch_1987 . In contrast, the non-equilibrium nature of PTCs allows exponentially growing modes in time, enabling light emission and amplification Lyubarov_2022 . This unique property makes temporally engineered band structures highly promising by enabling novel forms of spontaneous emission control Dikopoltsev_2022 ; Park_2025 , overcoming electromagnetic bounds inherent to equilibrium systems Hayran_2024 , and providing non-resonant and tunable gain, lasing, and frequency conversion with reduced phase-matching requirements compared to conventional nonlinear optical methods Lyubarov_2022 ; Khurgin_2024 .
However, experimental realization of a PTCs has also been recognized to be an arduous task Asgari_2024 ; Hayran_2022 ; Saha_2023 ; Hayran_2025 . Much like SPCs require a large dielectric contrast on wavelength scales to reshape the photonic band structure in a sizeable fashion, PTCs demand relative modulation depths of the optical properties to be of the order of unity (e.g. $\delta n/n\approx 1$ with $n$ the refractive index) while simultaneously providing ultrafast coherent modulations at optical cycle rates. From a material perspective Saha_2023 , dielectrics offer ultrafast response times but fail to produce such large modulations depths, while epsilon-near-zero materials produce larger depths but periodically and coherently driven modulations on optical-cycle timescales remains yet to be demonstrated Alam_2016 ; Khurgin_2020 . Furthermore, the estimated $\mathrm{TW/cm^{3}}$ power densities required to produce such nonlinearities are comparable with material damage thresholds, casting doubts on the feasibility Hayran_2022 ; Hayran_2025 . For these reasons, practical implementations of photonic time crystals have so far been ascribed to systems operating at electronic frequencies Reyes_Ayona_2015 ; Wang_2023 ; Lee2025 and the question remains opened as to their possible experimental realization at higher frequencies in an optical context.
Here, we show that this path is experimentally possible and demonstrate the first optical realization of a photonic time crystal. Using a multi-cycle Terahertz-frequency light field of modest amplitude ($\approx 40$kV/cm), we periodically drive a surface plasmonic cavity metamaterial and achieve strong (near-unity) and coherent (sub-optical-cycle) parametric modulation of its resonance, an effect which arises from dynamical variations in the carriers’ effective mass and kinetic energy. Our results provide the basis for time-crystalline phenomena in plasmonic systems Kiselev_2024 ; Feinberg_2025 ; Kiselev2024 , including the parametric amplification of plasmons and the generation of entangled plasmons, all within the experimentally accessible regime of parameters of this platform. Moreover, the combination of metamaterial design and plasmonic tunability Aupiais_2023 ; Aupiais_2024 offers a promising route to enrich PTCs with the full spectrum of photonic degrees of freedom.
Experiment
To demonstrate our approach, we take a conceptually simple metamaterial structure, namely a periodic lattice of sub-wavelength metal/insulator/plasma cavities, where the plasmonic material corresponds to the small band-gap semiconductor InSb Aupiais_2023 . It is shown in Fig. 1 together with the experimental configuration.
Figure 2: Temporal response of the metamaterial under a multicycle drive with a peak field of 40kV/cm: frequency-resolved power reflectivity (central panels) and phase (right panels) spectra as a function of time $\tau$ during the drive (left panels) a) Coarse grained dynamics measured over the complete duration of the drive b) Close up on the dynamics in the vicinity of the peak driving field with sub-cycle temporal resolution. c) Comparison of the sub-cycle dynamics shown in b) obtained from a parametrically driven cavity model (see text for details).
The equilibrium power reflectivity of the metamaterial, shown in Fig. 1b, exhibits an absorption dip centered at $0.77\rm{THz}$ with a full width at half maximum of $\sim 0.18\rm{THz}$. It corresponds to the resonance of the fundamental $\lambda/2$ mode of the sub-wavelength Fabry-Perot surface plasmonic cavity, whose mode profile is displayed in Fig. 1c. The periodically driven regime is investigated using a multi-cycle THz pump pulse centered at $0.69\rm{THz}$—slightly below the equilibrium resonance frequency—with a peak field amplitude of approximately 40kV/cm. The system’s dynamics during the drive is monitored using a broadband THz probe pulse, $\tilde{E}(\tau,t)$, where $\tau$ denotes the delay between the pump and the electro-optic sampling beam, and $t$ represents the internal delay between the sampling beam and the THz probe pulse, as defined in Fig.1a. For each time delay $\tau$, the frequency-resolved response of the metamaterial is obtained by performing a Fourier transform over the internal delay $t$.
Fig. 2a-b displays the temporal evolution of the power reflectivity and phase of the metamaterial’s response, as measured experimentally. Fig. 2a shows the metamaterial’s response across the full duration of the drive with coarse temporal resolution, while Fig.2b provides a close-up view in the vicinity of the peak of the driving field with sub-cycle temporal resolution. During the interaction, we observe a pronounced redshift of the cavity resonance, with an amplitude that closely tracks the envelope of the driving field. At large positive delays-when the driving field has almost vanished -the metamaterial response returns to its equilibrium state without any signature of a long-lived dynamics nor heating, suggesting a coherent interaction between the drive and the metamaterial. The measurement at sub-cycle temporal resolution in Fig. 2b fully confirms this picture, revealing pronounced and periodic modulations in both reflectivity and phase, which are driven on a timescale comparable to the optical cycle. Following an initial buildup of the response, the metamaterial reaches a quasi–steady-state characterized by 1) a red-shift of the metamaterial resonance frequency on average and 2) forced oscillations of the resonance parameters around this state. As can be seen from the dotted lines, these oscillations occur twice per optical cycle, corresponding to the second harmonic of the drive frequency. Below we model this system from basic principles to gain physical insights into its dynamics and quantify the effect of the time-modulation on the metamaterial optical properties.
Model
We attribute the temporal modulations of the surface plasmon cavity resonance to dynamical variations in the carriers’ effective mass $m^{*}(\tau)$ and kinetic inductance associated with non-parabolic conduction bands in small band-gap semiconductors Houver_2019 ; Soranzio_2024 : $m^{*}(\tau)=m^{*}_{0}+\delta m^{*}(\tau)$, where $m^{*}_{0}$ corresponds to the carriers’ rest effective mass. Under periodic driving at frequency $\omega_{\mathrm{d}}$, the electrons’ momentum $p(\tau)$ oscillates as $p(\tau)=p_{0}\cos(\omega_{\mathrm{d}}\tau)$. For a non-parabolic and isotropic conduction band given by the Kane model Kane_1957 , this results in effective mass modulations at twice the drive frequency:
$\delta m^{*}(\tau)=\alpha p^{2}(\tau)=\frac{\alpha}{2}{p_{0}}^{2}+\frac{\alpha}{2}{p_{0}}^{2}\cos(2\omega_{\mathrm{d}}t)$ (1)
where $\alpha$ accounts for the deviation from a purely parabolic band dispersion ($\alpha>0$ here, see methods). The resulting oscillation of the plasmonic cavity resonance can be qualitatively understood as follows. Using an LC-resonator model for the surface plasmonic cavity, the resonance frequency is given by $\omega_{\mathrm{c}}=\frac{1}{\sqrt{LC}},$ where $L$ and $C$ are the inductance and capacitance of the mode, respectively. For surface plasmons, the inductance $L=L_{G}+L_{K}$ includes contributions from both a geometric inductance $L_{G}$, arising from retardation effects, and a kinetic inductance $L_{K}\propto\frac{m^{*}}{ne^{2}}$, which represents part of the mode energy that is stored into the kinetic energy of the carriers (with $n$ the carrier density and $e$ the charge) Staffaroni_2012 . The kinetic inductance can be separated into an equilibrium and a time-varying part: $L_{K}=L_{K}^{0}+\delta L_{K}(\tau)$. Rigorously, it can be shown that $\delta L_{K}(\tau)/L_{K}^{0}=\beta\delta m^{*}(\tau)/m^{*}_{0}$ where the time-independent parameter $0\leq\beta\leq 1$ reaches unity only in the case of negligible geometric inductance and is lower otherwise (see methods and supplementary material). The resulting modulation of the resonance frequency is:
$\omega_{\mathrm{c}}(\tau)\approx\omega_{\mathrm{c}}^{0}\left(1-\frac{1}{2}\frac{L_{K}^{0}}{L_{G}+L_{K}^{0}}\beta\frac{\delta m^{*}(\tau)}{m^{*}_{0}}\right)$ (2)
where $\omega_{\mathrm{c}}^{0}=\frac{1}{\sqrt{(L_{G}+L_{K}^{0})C}}$ is the equilibrium resonance frequency. This equation has a simple physical interpretation: since modulation of the cavity frequency originates from variations of the kinetic inductance via the carriers’ effective mass, its depth is controlled by the relative contributions of the kinetic and geometric inductance to the cavity mode. The presence of a geometric inductance $L_{G}$, which is unaffected by the drive, acts only to reduce the overall depth of the cavity modulation. Taken together, equations (1) and (2) predict the red-shift of the cavity resonance and its modulation around the shifted resonance at the second harmonic of the drive frequency, as observed experimentally.
Such qualitative description can be put on solid grounds starting from the Hamiltonian and canonical quantization procedures. Specifically, excitations of the surface plasmonic cavity field described by their creation and annihilation operators $\hat{a}$ and $\hat{a}^{\dagger}$ evolve according to the following Hamiltonian:
$\hat{H(\tau)}=\hbar\omega_{\rm c}^{0}\hat{a}^{\dagger}\hat{a}-\frac{\hbar\eta(\tau)}{2}\left(\hat{a}^{\dagger}-\hat{a}\right)^{2}$ (3)
where the time-dependent modulation parameter is given by $\eta(\tau)=-\frac{\omega_{\rm c}^{0}}{2}\frac{L_{K}^{0}}{L_{G}+L_{K}^{0}}\beta\frac{\delta m^{*}(\tau)}{m^{*}_{0}}$ (see methods and supplementary material). This Hamiltonian describes a parametrically driven cavity resonance and the emergence of an interaction between surface plasmons that is generated by the drive. To capture the metamaterial response function that is measured experimentally -the reflectivity-, we implement an input-output formalism with two baths representing radiative and non-radiative (ohmic) decay channels. These account for coupling to an external electromagnetic probe and material losses, respectively. Adopting the experimental temporal conventions and using a single adjustable parameter, we present in Fig. 2c the computed time-resolved reflection coefficient $r(\tau,\nu)$ that best matches the experimental observations. This approach allows direct comparison between experimental and theoretical response functions, enabling inference of the time-dependent modulation of microscopic system parameters.
Discussion
We obtain a maximum modulation parameter $\eta_{0}/\omega_{\mathrm{c}}^{0}=\max\limits_{\tau}|\eta(\tau)|/\omega_{\mathrm{c}}^{0}=0.18\pm 0.02$ for which the model reproduces the experimental data with very good accuracy. This corresponds to an impressive $\delta m^{*}(\tau)/m^{*}_{0}\approx 80\%$ dynamical effective mass modulations. As a function of driving strength, we further establish that these parametric effects are saturated for the reported peak driving field of 40kV/cm. This is indicated by the leveling-off of the response and the emergence of incoherent nonlinearities measured at a larger field of $140\textrm{kV/cm}$ (see supplementary material). Hence, this value appears to be a physical bound for coherent effective mass modulations in this system. Given this value, a strategy to enhance further parametric effects consists in increasing the ratio $L_{K}^{0}/L_{G}$. It can be achieved by decreasing the size of the plasmonic cavities so as to operate in the ’electrostatic’ regime where retardation effects vanish and kinetic ones dominate Staffaroni_2012 ; Greffet_2012 : $L_{K}^{0}\gg L_{G}$ and $\beta=1$. In the following, using Floquet analysis with periodic drive $\eta(\tau)=-\eta_{0}\cos^{2}(\omega_{\mathrm{d}}\tau)$ and accounting for the observed losses, we highlight three key physical phenomena lying within the experimentally accessible range of parameters of this system for : 1) the present cavity design ($\beta\approx 0.65,L_{K}^{0}/L_{G}\approx 2.4$, $\eta_{0}/\omega_{\mathrm{c}}^{0}\approx 0.18$) and 2) an electrostatic plasmonic cavity ($\beta=1,L_{K}^{0}\gg L_{G}$, $\eta_{0}/\omega_{\mathrm{c}}^{0}=0.4$).
Parametric amplification of surface plasmons
As well known from other parametrically driven systems, the modulation induced by the drive generates gain on the cavity mode. The corresponding growth rate of the mode $\mathrm{Im}(\mu)/\omega_{\mathrm{c}}^{0}$, with $\mu$ being an eigenvalue of the Floquet matrix describing long-term behavior of the system (see methods), is reported in the phase diagram of Fig. 3a as a function of drive frequency and modulation strength, together with the operating points of the two aforementioned cavity designs. When the growth rate overcomes the total loss rate $\Gamma/\omega_{\mathrm{c}}^{0}$ (including radiative and non-radiative contributions), the system develops a parametric instability: it crosses a threshold illustrated by the solid contour line and defined by $\mathrm{Im}(\mu)/\omega_{\mathrm{c}}^{0}=\Gamma/\omega_{\mathrm{c}}^{0}$ in Fig. 3a for which quantities diverge exponentially and parametric amplification sets in. Even before reaching this threshold and within the broader region defined by the dashed contour line, the system remains stable as a function of time but demonstrates average reflectivities above 1 (see methods and supplementary material). This highlights the strong potential of this platform for achieving light amplification and lasing of surface plasmons.
Figure 3: Plasmonic time crystal phenomena computed for the experimentally accessible regime of parameters of this platform
a) Stability diagram: growth rate of the cavity mode as a function of drive frequency $\omega_{\mathrm{d}}$ and modulation strength $\eta_{0}$ relative to the equilibrium cavity frequency $\omega_{\mathrm{c}}^{0}$. The solid contour line defined by $\mathrm{Im}(\mu)/\omega_{\mathrm{c}}^{0}=\Gamma/\omega_{\mathrm{c}}^{0}=0.12$ marks the threshold for parametric amplification given the observed losses. Dashed contour line indicate regions exhibiting average reflectivity above one. Operating points of the present cavity design (star) and an electrostatic cavity (dot) are indicated (see text) b) Wigner function of the cavity state accounting for the losses. Top panel: no drive. Middle panel: this experiment’s parameters ($\eta_{0}/\omega_{\mathrm{c}}^{0}=0.18$, $\omega_{\rm d}/\omega_{\mathrm{c}}^{0}=0.92$) Bottom panel: an electrostatic cavity ($\eta_{0}/\omega_{\mathrm{c}}^{0}=0.4$, $\omega_{\rm d}/\omega_{\mathrm{c}}^{0}=0.8$) c Plasmonic time crystal $k$-gap openings. Real and imaginary parts of the eigenmodes’ frequency $\omega$ as a function of momentum (or Bloch-momentum) $k$, normalized to the plasma frequency $\omega_{\mathrm{p}}$ and including the losses, for the equilibrium (dashed lines) and the temporally-modulated systems (solid lines). Top panel: a dispersive insulator/plasmonic interface. Bottom panel: our dispersion-free plasmonic metamaterial design for cavities tuned in the electrostatic regime (shown in the first Brillouin zone of the metamaterial of period $d$. The drive frequency corresponds to $\omega_{\mathrm{d}}/\omega_{\mathrm{p}}=0.68$ in both cases.
Generation of squeezed and entangled plasmons
[…]
$k$-agnostic gap openings in plasmonic metamaterial time crystals
Moving from a single cavity picture to the actual periodically structured metamaterial, a central aspect of time crystals lies in the opening of $k$-gaps in their dispersion relations Asgari_2024 , where $k$ denotes the momentum (translational invariant systems) or quasi-momentum (periodically structured metamaterials) of photonic eigenmodes. Since material dispersion limits the width of the $k$-gaps, several strategies were recently proposed to widen their span Wang_2024 ; Feinberg_2025 . The ability of our metamaterial approach to engineer a flat dispersion allows to address such challenges Aupiais_2023 . In Fig 3c. we compare a dispersive insulator/plasmonic interface together with our metamaterial design and compute $k$-gap openings in the dispersion relation (see methods). For the dispersive plasmonic interface (top panel), although the $k$-gaps are already significant, they fail to cover the whole dispersion. In contrast, our dispersion-free metamaterial (bottom panel) achieves $k$-gap openings across the full Brillouin zone, effectively realizing a $k$-agnostic time crystal.
Conclusion
In conclusion, we report the first optical implementation of a photonic time crystal. This is realized in a plasmonic system and enabled by the sizeable and coherent dynamical variations of the carriers’ effective mass. Our work establishes that this mechanism, previously known in bulk systems, can be extended to surface plasmons and leveraged as a foundational element for constructing plasmonic metamaterials harbouring Floquet physics and time crystalline phenomena. Along those lines, the inherent tunability of the plasmonic medium to external parameters Aupiais_2023 ; Aupiais_2024 offers a unique route towards extension of PTC phenomena to other photonic degrees of freedom, including spin and orbital ones.
Acknowledgments
Y.L. acknowledges support from Agence Nationale de la Recherche (grant n ANR-23-CE30-0030) and Region Ile-de-France in the framework of DIM QuanTiP. M.S. acknowledges financial support from the ERC consolidator Grant No. 101002955 - CONQUER. T.G. acknowledges support from Chinese Scholarship Council (CSC202206210114). G.M.A. acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101146870 – COMPASS. Parts of this research were carried out at ELBE at the Helmholtz-Zentrum Dresden-Rossendorf e.V., a member of the Helmholtz Association.
Methods
Sample manufacturing
A $\langle 100\rangle$-oriented, nominally undoped InSb crystal with dimensions of $5\penalty 10000\ \mathrm{mm}\times 5\penalty 10000\ \mathrm{mm}\times 0.5\penalty 10000\ \mathrm{mm}$ is used as the starting material. A $\sim 3\penalty 10000\ \mu\mathrm{m}$-thick $\mathrm{Si}_{3}\mathrm{N}_{4}$ insulating layer is first deposited via plasma-enhanced chemical vapor deposition. The top metallic pattern is then defined through photolithography, followed by Ti/Au (20 nm/200 nm) evaporation and lift-off. The resulting metamaterial structure consists of periodic metallic stripes with a width of $s=41\penalty 10000\ \mu\mathrm{m}$ and spacing $a=16\penalty 10000\ \mu\mathrm{m}$, yielding a total period of $d=s+a=57\penalty 10000\ \mu\mathrm{m}$ and follows the design reported in a previous work Aupiais_2023 .
THz pump-probe spectroscopy
The driving beam is generated by the superradiant Terahertz source TELBE at the ELBE accelerator (Helmholtz-Zentrum Dresden-Rossendorf). It produces tunable narrow-band and multi-cycle pulses from $0.1$ to $3$ THz and a selectable repetition rate from 10 to 250 kHz. For this experiment, the source was tuned at a central frequency of $0.69\penalty 10000\ \mathrm{THz}$ with a $50\penalty 10000\ \mathrm{kHz}$ repetition rate. The broadband THz probe is generated by a photo-conductive emitter (LaTera 10/10, HZDR Innovation) driven by a $100\penalty 10000\ \mathrm{kHz}$ ultrafast Ti:Sapphire amplifier system synchronized to TELBE via a stabilized fiber link. THz pulse measurements are achieved via a balanced detection and electro-optical sampling in a $100\penalty 10000\ \mu\mathrm{m}$-thick $\langle 110\rangle$ ZnTe crystal. THz pump and probe, combined with a HRFZ Silicon plate, are co-propagating and co-polarized and are focused onto the sample at $45^{\circ}$ incidence angle with s-polarization in a spot of FWHM diameter around $900\penalty 10000\ \mu\mathrm{m}$. All measurements are performed at room temperature in a purged atmosphere.
Time-domain windowing is applied to isolate the main THz pulse from other time-delayed reflections and is Fourier transformed to obtain power reflectivity and phase spectra. The reflection coefficient is obtained based on a self-referencing method by normalizing the TM response to the TE response of the sample: TM-polarized THz fields (perpendicular to the stripes) excite cavity modes, while TE-polarized fields (parallel to the stripes) are fully reflected and are used as a reference. To avoid nonphysical frequency-independent contributions to the signal, as can arise from variations in the generation or balancing conditions of the detection, power reflectivity spectra are scaled to ensure reaching the same limit in the low frequency part of the spectra ($0.2$-$0.45\penalty 10000\ \mathrm{THz}$).
Theoretical model
The collective motion of electrons leading to the surface plasmon cavity resonance is described as a $LC$ oscillator. The Hamiltonian can be expressed as:
$H=\frac{1}{2C}Q^{2}+\frac{1}{2}LI^{2}$ (4)
Here, $Q$ and $I$ are collective variables describing the dynamics of the surface plasmon cavity field. They are associated respectively with the charge displacement and the plasmonic current and are related by charge conservation ($\dot{Q}=I$). $C$ and $L$ parametrize the capacitive and inductive parts of the energy of the plasmon field.
The inductance $L$ has two contributions Staffaroni_2012 :
  • a geometric inductance $L_{G}$ arising from retardation effects of the electromagnetic field
  • a kinetic inductance $L_{K}\propto\frac{m^{*}}{ne^{2}}$ characterizing energy that is stored into the kinetic energy of the carriers ($m^{*}$,$n$ and $e$ are the mass, density and charge of the carriers).
    We decompose it as a sum of the rest kinetic inductance of the carriers $L_{K}^{0}$ and a time-dependent contribution $\delta L_{K}(\tau)$ arising from effective mass modulations: $L_{K}=L_{K}^{0}+\delta L_{K}(\tau)$.
The total inductance is the sum of these different contributions:
$L=L_{G}+L_{K}^{0}+\delta L_{K}(t)$ (5)
With these definitions, the Hamiltonian writes :
$H(\tau)=\frac{1}{2C}Q^{2}+\frac{1}{2}(L_{G}+L_{K}^{0})I^{2}+\frac{1}{2}\delta L_{K}(\tau)I^{2}$ (6)
After quantization, this can be expressed with ladder operators $\hat{a}$ and $\hat{a}^{\dagger}$ describing excitations of the surface plasmon cavity field as:
$\hat{H(\tau)}=\hbar\omega_{\rm c}^{0}\hat{a}^{\dagger}\hat{a}-\frac{\hbar\eta(\tau)}{2}\left(\hat{a}^{\dagger}-\hat{a}\right)^{2}$ (7)
where the modulation parameter $\eta(\tau)$ is:
$\eta(\tau)=-\frac{\omega_{\rm c}^{0}}{2}\frac{\delta L_{K}(\tau)}{L_{G}+L_{K}^{0}}$ (8)
The modulation parameter can be also expressed as a function of the effective mass modulations as:
$\eta(\tau)=-\frac{\omega_{\rm c}^{0}}{2}\frac{L_{K}^{0}}{L_{G}+L_{K}^{0}}\beta\frac{\delta m(\tau)}{m^{*}_{0}}$ (9)
where $0\leq\beta\leq 1$ is a time-independent factor that links the variations of the kinetic inductance to the variations of the effective mass: $\frac{\delta L_{K}(\tau)}{L_{K}^{0}}=\beta\frac{\delta m^{*}(\tau)}{m^{*}_{0}}$. Its computation is outlined below. Equation (7) is supplemented with an input-ouput formalism to compute the reflection coefficient, as detailed in the supplementary material.
Model parameters
Computation of the parameters entering equations (8) and (9) is outlined here and explained in more detail in the supplementary material. These calculations are based on the determination of the cavity resonance frequency $\omega_{\rm c}$, obtained by solving the dispersion relation of a truncated metal/insulator/plasmonic waveguide Aupiais_2023 , when parameters are varied. Notably, the resonance frequency depends, among other geometrical and optical parameters of the constitutive elements of the metamaterial, on the plasma frequency of the plasmonic medium supporting the resonance: $\omega_{\rm p}=\sqrt{\frac{ne^{2}}{\epsilon_{0}\epsilon_{\infty}m^{*}}}$, where $n$ and $e$ are the density and charge of the carriers, $m^{*}$ their effective mass, $\epsilon_{0}$ the vacuum permittivity and $\epsilon_{\infty}$ the high-frequency permittivity of the plasmonic medium. Therefore, the ratio of kinetic to geometric inductance $L_{K}^{0}/L_{G}$ at equilibrium can be determined from the resonance frequencies $\omega_{\mathrm{c}}^{0}$ and $\omega_{0}$ obtained by varying the plasma frequency for two otherwise identical cavities: 1) with equilibrium plasma frequency, yielding $\omega_{\rm c}=\frac{1}{\sqrt{(L_{K}^{0}+L_{G})C}}$ and 2) with infinite plasma frequency, for which the kinetic inductance vanishes while the geometric inductance and capacitance remain unchanged, yielding $\omega_{0}=\frac{1}{\sqrt{L_{G}C}}$. Similarly, the dependence of kinetic inductance with the effective mass, allowing to determine $\delta L_{K}$ and $\beta$, is obtained through a variation of the effective mass that enters the plasma frequency, all other parameters being fixed. We point out that these computations are parameter-free and obtained from experimental measurements of the geometrical and optical properties of the constitutive elements of the metamaterial, as detailed in Aupiais_2023 .
Temporal dynamics (Figure 2)
To compute the temporal response of the system to the experimentally measured incident driving field $E_{d}(\tau)$ shown in Fig. 2, we determine $\eta(\tau)$ from equation (9) as follows. We consider a non-parabolic band structure given by the Kane model suitable for the description of small-band gap semiconductors Kane_1957 where the electronic dispersion relation $\varepsilon(p)$ satisfies $\varepsilon(p)(1+\alpha\varepsilon(p))=\frac{p^{2}}{2m^{*}_{0}}$ and $\alpha$ is the parameter that accounts for non-parabolicity. The effective mass $m^{*}(p)$ is defined by $m^{*}(p)=p\left(\frac{d\varepsilon}{dp}\right)^{-1}$, which allows to express it as: $m^{*}(p)=m^{*}_{0}(1+2\alpha\varepsilon(p))$ Zawadzki_1974 . Keeping only the leading order term $\propto p^{2}$ in the electronic dispersion relation, we parametrize the effective mass as $m^{*}(p)=m^{*}_{0}(1+2\alpha\frac{p^{2}}{2m^{*}_{0}})$, so that $\frac{\delta m^{*}(p)}{m^{*}_{0}}=\frac{\alpha p^{2}}{m^{*}_{0}}$. We solve the dynamics of the electrons’ momentum $p(\tau)$ in the resonator according to:
$\frac{d^{2}p}{d\tau^{2}}+2\Gamma\frac{dp}{d\tau}+{\omega_{\mathrm{c}}^{0}}^{2}p$ $=-e\frac{d}{d\tau}(fE_{\mathrm{d}}(\tau))$ (10)
where the resonator frequency is kept fixed for simplicity and where $f$ accounts for the fraction of the incident field $fE_{\mathrm{d}}(\tau)$ that effectively enters the plasmonic medium and couples to the electrons’ momentum. It is the only free parameter of our model. Solving this equation for $p(\tau)$, we determine $\eta(\tau)=-\frac{\omega_{\rm c}^{0}}{2}\frac{L_{K}^{0}}{L_{G}+L_{K}^{0}}\beta\frac{\alpha p(\tau)^{2}}{m^{*}_{0}}$ and compute the quantum dynamics eq. (7) with an input-output formalism to obtain the reflection coefficient plotted in Fig. 2c.
The parameters used in the computation are listed in Table 1.
Table 1: Model Parameters
Parameter Name Notation Value (SI Unit) Source
Non-parabolicity parameter $\alpha$ $5.6\penalty 10000\ \mathrm{eV}^{-1}$ Zawadzki_1974
Radiative power decay rate $\gamma$ $2\pi\times 0.042\penalty 10000\ \mathrm{THz}$ Experimental fit
Non-radiative power decay rate $\gamma^{\prime}$ $2\pi\times 0.141\penalty 10000\ \mathrm{THz}$ Experimental fit
Total field decay rate $\Gamma=(\gamma+\gamma^{\prime})/2$ $2\pi\times 0.091\penalty 10000\ \mathrm{THz}$ Experimental fit
Cavity resonance frequency $\omega_{\mathrm{c}}^{0}$ $2\pi\times 0.77\penalty 10000\ \mathrm{THz}$ Experimental fit
Effective drive field $f$ $27\%$ Free parameter
Inductance ratio $L_{K}^{0}/L_{G}$ $2.4$ Computed
Inductance to effective mass variations $\beta$ $0.65$ Computed
Floquet analysis (Figure 3)
For the Floquet analysis presented in Fig. 3, we use a periodic drive: $\eta(\tau)=-\eta_{0}\cos^{2}(\omega_{\mathrm{d}}\tau)$.
Phase diagram (Figure 3a)
Under a periodic drive $\eta(t)$, the evolution operator of the cavity mode (a^,a^)($\hat$$^{\dagger},\hat{a}$) is decomposed using Floquet’s theorem as:
$\mathbf{U}(\tau)=\mathbf{G}(\tau)e^{-\mathrm{i}\mathbf{\mu}\tau},$ (11)
where $\mathbf{G}(\tau)$ and $\mathbf{\mu}$ are both $2\times 2$ matrices acting on the spinor space of the cavity mode (a^,a^)($\hat$$^{\dagger},\hat{a}$) (see supplementary material). $\mathbf{G}(\tau)$ is periodic with the same period as the drive and describes the periodic micro-motion of the system while the Floquet exponent matrix $e^{-\mathrm{i}\mathbf{\mu}t}$ describes its long-term behavior. In the absence of dissipation, the imaginary part of the eigenvalues $\mu_{1,2}$ of matrix $\mathbf{\mu}$ determine wether the system is stable ($\mathrm{Im}(\mu_{1,2})<0$) or unstable ($\mathrm{Im}(\mu_{1,2})>0$), as quantities evolve as $e^{\mathrm{Im}(\mu_{1,2})t}$ at long times. In the presence of dissipation characterized by a total decay rate $\Gamma=\frac{\gamma+\gamma^{\prime}}{2}$, it can be shown that eigenvalues of $\mathbf{\mu}$ undergo a shift by $\Gamma$ while $\mathbf{G}(t)$ remains unaffected. Defining $\mathrm{Im}(\mu)=\max\limits_{1,2}\mathrm{Im}(\mu_{1,2})$, this allows to conclude directly on the stability of the dynamics from a contour line defined by $\frac{|\mathrm{Im}(\mu)|}{\omega_{\mathrm{c}}^{0}}=\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$ and which marks the transition between stable ($\frac{|\mathrm{Im}(\mu)|}{\omega_{\mathrm{c}}^{0}}<\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$) and unstable dynamics ($\frac{|\mathrm{Im}(\mu)|}{\omega_{\mathrm{c}}^{0}}>\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$). The total normalized decay rate $\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}=0.12$ presented in Fig. 3a. corresponds to the experimentally measured one and can be found in Table 1.
Wigner function (Figure 3b)
The Wigner function distribution of the cavity state, plotted in Fig. 3b., is defined via its density matrix $\hat{\rho}(\tau)$ as :
$W(\alpha,\tau)=\frac{2}{\pi}\text{Tr}\left(\hat{\rho}(\tau)\hat{D}(\alpha)(-1)^{a^{\dagger}a}\hat{D}(-\alpha)\right),$ (12)
where for all $\alpha\in\mathbb{C}$, the displacement operator is defined by $\hat{D}(\alpha)=\exp(\alpha a^{\dagger}-\alpha^{\ast}a)$. It is computed from the time evolution of the density matrix according to its Hamiltonian eq.(7) and accounting for losses at the rate $\Gamma=\frac{\gamma+\gamma^{\prime}}{2}$ as :
$\partial_{t}\hat{\rho}=-\mathrm{i}\left[\hat{H}(\tau),\hat{\rho}\right]+\frac{\gamma+\gamma^{\prime}}{2}\mathcal{D}[a]\hat{\rho}$. (13)
where $\mathcal{D}[a]$ is the the one-photon loss dissipator which acts on the density matrix through $\mathcal{D}[a]\rho=a\rho a^{\dagger}-\frac{1}{2}\left\{a^{\dagger}a,\rho\right\}$. Further details of the calculations are given in the supplementary materials.
Photonic time crystal (Figure 3c)
To treat the case of photonic time crystals, we extend the formulation of the model to general dispersive systems. Eigenmodes depend on index $k$, which corresponds either to the parallel wavevector (translationally-invariant systems) or Bloch-wavevector (periodically structured metamaterials) of a given eigenmode. For such systems, the Hamiltonian writes (see supplementary materials):
$\hat{H(\tau)}=\sum_{k}\hbar(\omega_{\rm c}^{0}(k)+\eta_{k}(\tau))\hat{a}_{k}^{\dagger}\hat{a}_{k}-\frac{\hbar\eta_{k}(\tau)}{2}\left(\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}+{\hat{a}}_{k}{\hat{a}}_{-k}\right)$ (14)
where $\omega_{\rm c}^{0}(k)$ and $\eta_{k}(\tau)$ are, respectively, the $k$-dependent equilibrium resonance frequency and modulation parameter of the eigenmodes. As we consider systems that are driven homogeneously at $k=0$, creation and annihilation of plasmon pairs have opposite momenta in order to satisfy momentum conservation rules.
Computation of $\omega_{\rm c}^{0}(k)$ and $\eta_{k}(\tau)$ is performed with an identical procedure as described above for a single cavity: the dispersion relation of the system is computed both at equilibrium and through variation of the effective mass that enters the plasma frequency, allowing for the determination of the relation of proportionality as in eq. 9 between $\frac{\eta_{k}(\tau)}{\omega_{\rm c}^{0}(k)}$ and $\frac{\delta m^{*}(\tau)}{m_{0}^{*}}$ for any given $k$-mode. Floquet analysis is then performed on Hamiltonian eq. 14 in order to determine the Floquet eigenvalues $\omega_{1,2}$ of the driven system and to track the opening of $k$-gaps. After accounting for non-radiative losses with a shift in the imaginary part and normalizing by the equilibrium plasma frequency $\omega_{\mathrm{p}}$, real and imaginary parts are reported in Fig. 3c in the two following cases for: 1) a dispersive insulator/plasmonic interface and 2) our dispersion-free metamaterial design for cavities tuned in the electrostatic regime.

### v2

@@
 Abstract
-Periodically driven optical materials and metamaterials have recently emerged as a promising platform for realizing photonic time crystals (PTCs)—systems whose optical properties are strongly and periodically modulated on time scales comparable to the optical cycle of light Asgari_2024 .
+Periodically driven optical materials and metamaterials have recently emerged as a promising platform for realizing photonic time crystals (PTCs), which are systems whose optical properties are strongly and periodically modulated on timescales comparable to the optical cycle of light Asgari et al.
+(2024).
 These time-varying structures are the temporal counterparts of spatial photonic crystals (SPCs), for which a large and periodic dielectric contrast is achieved spatially on wavelength scales.
-Just as SPCs have revolutionized control over light–matter interactions by engineering the photonic density of states in space Yablonovitch_1987 ; John_1987 ; Joannopoulos_2011 ; Lu_2014 , PTCs promise comparable breakthroughs from a fundamentally new perspective: a temporal one Lustig_2018 ; Lyubarov_2022 ; Dikopoltsev_2022 ; Li_2023 ; Wang_2024 ; Park_2025 .
-However, harnessing such phenomena at optical frequencies poses severe experimental challenges Asgari_2024 ; Hayran_2022 ; Saha_2023 ; Hayran_2025 , as it requires order-unity modulation depths of the optical properties at optical cycle rates, a regime that has remained elusive to date.
-Here, we report the first optical realization of a photonic time crystal, achieved with a surface plasmon cavity metamaterial operating at Terahertz frequencies.
-We demonstrate strong (near-unity) and coherent (sub-optical cycle) periodic driving of the plasmonic metamaterial enabled by field-induced dynamical modulation of the carriers’ kinetic energy and effective mass—reaching up to 80% of their rest mass, an exceptionally high value that forms the basis for time crystalline phenomena with plasmons Kiselev_2024 ; Feinberg_2025 .
-Our experimentally informed theory reveals rich physics within the experimentally accessible parameter regime of this system, including parametric amplification and entangled plasmon generation, and establishes a robust new platform for time-domain photonics.
+Just as SPCs have revolutionized control over light–matter interactions by engineering the photonic density of states in space Yablonovitch (1987); John (1987); Joannopoulos et al.
+(2011); Lu et al.
+(2014), PTCs promise comparable breakthroughs from a fundamentally new perspective: a temporal one Lustig et al.
+(2018); Lyubarov et al.
+(2022); Dikopoltsev et al.
+(2022); Li et al.
+(2023); Wang et al.
+(2024); Park et al.
+(2025).
+However, harnessing such phenomena all-optically poses severe experimental challenges Asgari et al.
+(2024); Khurgin et al.
+(2020); Hayran et al.
+(2022); Saha et al.
+(2023); Hayran and Monticone (2025), as it requires order-unity modulation depths of a material’s optical properties on ultrafast timescales comparable to the light cycle, a regime that has remained elusive to date.
+Here, we demonstrate the first all-optical realization of a photonic time crystal, achieved with a surface plasmon cavity metamaterial operating at Terahertz (THz) frequencies.
+We demonstrate strong (near-unity) and coherent (sub-optical cycle) periodic driving of the plasmonic metamaterial enabled by field-induced dynamical modulation of the carriers’ kinetic energy and effective mass—reaching up to 80% of their rest mass.
+Our spectroscopic measurements reveal a transition into the PTC regime mediated by an exceptional point, at which two Floquet-driven optical eigenmodes coalesce.
+In the PTC regime, emergent gain is shown to reduce plasmonic losses by more than 50% Kiselev et al.
+(2024b); Feinberg et al.
+(2025) and we predict plasmonic lasing to be within experimental reach.
+These results pave the way for temporal engineering of losses and light-matter interactions in plasmonic systems, and establish a robust new platform for time-domain photonics.
 Introduction
 […]
-Systems subjected to periodic temporal modulations, known as Floquet systems, have been extensively studied across diverse platforms, including mechanical Trainiti_2016 , atomic Eckardt_2017 , and solid-state Wang_2013 ; Nova_2016 ; Cavalleri_2017 ; McIver_2019 ; Oka_2019 ; Rudner_2020 .
-In such systems, periodic driving is known to provide fascinating properties not accessible in equilibrium Bukov_2015 and the ability to tailor emergent properties through Floquet engineering of effective Hamiltonians Goldman_2014 ; Weitenberg_2021 .
-Over the past few years, these ideas have increasingly influenced the area of photonics.
-Periodic temporal modulation of optical systems enables time, in addition to space, to be utilized as a synthetic dimension for the tailoring of optical eigenmodes and response functions, an approach described within the framework of four-dimensional or Floquet metamaterials Engheta_2020 ; Yin_2022 ; Galiffi_2022 ; Sapienza_2023 ; Engheta_2023 ; Horsley_2023 ; Sisler_2024 .
-For spatially homogeneous yet ultrafast temporal modulations approaching the timescale of the optical cycle, the related concept of photonic time crystals (PTCs) has recently emerged Asgari_2024 ; Lustig_2018 ; Lyubarov_2022 ; Dikopoltsev_2022 ; Li_2023 ; Wang_2024 ; Park_2025 ; Kiselev_2024 ; Feinberg_2025 ; Pan_2023 ; Hayran_2024 ; Khurgin_2024 .
-As depicted in Figure 1a, PTCs can be viewed as the temporal analogues of spatial photonic crystals (SPCs) Yablonovitch_1987 ; John_1987 ; Joannopoulos_2011 ; Lu_2014 by permuting space and time coordinates on one hand and momentum and frequency on the other.
-In SPCs, spatial periodicity generates band replicas along the momentum axis and opens frequency gaps ($\omega$-gaps), whereas in PTCs, temporal periodicity produces replicas along the frequency axis and opens momentum gaps ($k$-gaps).
+Dynamic control of optical systems in time via periodic modulation, or Floquet driving, has recently emerged as a powerful paradigm to photonic engineering.
+For a modulation frequency $\omega_{d}$ that is much smaller than the light frequency of interest $\omega$ $(\omega_{d}\ll\omega)$, this approach has been used to engineer a discrete synthetic dimension from the sidebands ladder at $\omega\pm n\omega_{d}$ ($n$ being an integer), a concept formalized within the framework of four-dimensional or Floquet metamaterials Engheta (2020); Yin et al.
+(2022); Galiffi et al.
+(2022); Sapienza et al.
+(2023); Engheta (2023); Horsley and Pendry (2023); Sisler et al.
+(2024).
+When the modulation and light frequencies become commensurate together $(\omega_{d}\approx 2\omega)$, qualitatively new phenomena emerge that enable photonic band structures to be engineered dynamically through a temporal approach, an aspect which is captured by the recently introduced concept of photonic time crystals (PTCs) Asgari et al.
+(2024); Lustig et al.
+(2018); Lyubarov et al.
+(2022); Dikopoltsev et al.
+(2022); Li et al.
+(2023); Wang et al.
+(2024); Park et al.
+(2025); Kiselev et al.
+(2024b); Feinberg et al.
+(2025); Pan et al.
+(2023); Hayran and Monticone (2024); Khurgin (2024).
+It should be noted that PTCs are periodically driven optical systems that do not break the discrete time-translational invariance imposed by the drive and should thus not be confused with many-body time crystals whose physics originate specifically from the breaking of time-translational invariance Wilczek (2012); Sacha and Zakrzewski (2017).
+As depicted in Figure 1a, PTCs can be viewed as the temporal analogues of spatial photonic crystals (SPCs) Yablonovitch (1987); John (1987); Joannopoulos et al.
+(2011); Lu et al.
+(2014) by permuting space and time coordinates on one hand and momentum and frequency on the other.
+In SPCs (resp.
+PTCs), spatial (resp.
+temporal) periodicity generates band replicas along the momentum (resp.
+frequency) axis and whose hybridization with the original band opens frequency (resp.
+momentum) gaps ($\omega$-gaps, resp.
+$k$-gaps) near their crossings.
 Wave-behavior within or in the vicinity of these gaps is key in governing light–matter interactions and differ fundamentally between the two cases.
@@
 1a).
-In SPCs, only the decaying solution is physically allowed, a consequence of their time-invariant and energy-conserving nature, leading to inhibition of spontaneous emission Yablonovitch_1987 .
-In contrast, the non-equilibrium nature of PTCs allows exponentially growing modes in time, enabling light emission and amplification Lyubarov_2022 .
-This unique property makes temporally engineered band structures highly promising by enabling novel forms of spontaneous emission control Dikopoltsev_2022 ; Park_2025 , overcoming electromagnetic bounds inherent to equilibrium systems Hayran_2024 , and providing non-resonant and tunable gain, lasing, and frequency conversion with reduced phase-matching requirements compared to conventional nonlinear optical methods Lyubarov_2022 ; Khurgin_2024 .
-However, experimental realization of a PTCs has also been recognized to be an arduous task Asgari_2024 ; Hayran_2022 ; Saha_2023 ; Hayran_2025 .
-Much like SPCs require a large dielectric contrast on wavelength scales to reshape the photonic band structure in a sizeable fashion, PTCs demand relative modulation depths of the optical properties to be of the order of unity (e.g.
-$\delta n/n\approx 1$ with $n$ the refractive index) while simultaneously providing ultrafast coherent modulations at optical cycle rates.
-From a material perspective Saha_2023 , dielectrics offer ultrafast response times but fail to produce such large modulations depths, while epsilon-near-zero materials produce larger depths but periodically and coherently driven modulations on optical-cycle timescales remains yet to be demonstrated Alam_2016 ; Khurgin_2020 .
-Furthermore, the estimated $\mathrm{TW/cm^{3}}$ power densities required to produce such nonlinearities are comparable with material damage thresholds, casting doubts on the feasibility Hayran_2022 ; Hayran_2025 .
-For these reasons, practical implementations of photonic time crystals have so far been ascribed to systems operating at electronic frequencies Reyes_Ayona_2015 ; Wang_2023 ; Lee2025 and the question remains opened as to their possible experimental realization at higher frequencies in an optical context.
-Here, we show that this path is experimentally possible and demonstrate the first optical realization of a photonic time crystal.
-Using a multi-cycle Terahertz-frequency light field of modest amplitude ($\approx 40$kV/cm), we periodically drive a surface plasmonic cavity metamaterial and achieve strong (near-unity) and coherent (sub-optical-cycle) parametric modulation of its resonance, an effect which arises from dynamical variations in the carriers’ effective mass and kinetic energy.
-Our results provide the basis for time-crystalline phenomena in plasmonic systems Kiselev_2024 ; Feinberg_2025 ; Kiselev2024 , including the parametric amplification of plasmons and the generation of entangled plasmons, all within the experimentally accessible regime of parameters of this platform.
-Moreover, the combination of metamaterial design and plasmonic tunability Aupiais_2023 ; Aupiais_2024 offers a promising route to enrich PTCs with the full spectrum of photonic degrees of freedom.
+In SPCs, only the decaying solution is physically allowed, a consequence of their time-invariant and energy-conserving nature, leading to inhibition of spontaneous emission Yablonovitch (1987).
+In contrast, the non-equilibrium nature of PTCs allows exponentially growing modes in time, enabling light emission and amplification Lyubarov et al.
+(2022).
+This unique property makes temporally engineered band structures highly promising for novel forms of spontaneous emission control Dikopoltsev et al.
+(2022); Park et al.
+(2025), overcoming electromagnetic bounds inherent to equilibrium systems Hayran and Monticone (2024), and providing non-resonant and tunable gain, lasing, and frequency conversion with reduced phase-matching requirements compared to conventional nonlinear optical methods Lyubarov et al.
+(2022); Khurgin (2024).
+However, all-optical realizations of PTCs have proven to be a challenging task Asgari et al.
+(2024); Khurgin et al.
+(2020); Hayran et al.
+(2022); Saha et al.
+(2023); Hayran and Monticone (2025).
+Much like SPCs, which require a large dielectric contrast on wavelength scales to engineer photonic band structures, PTCs demand order-unity relative modulation depths of a material’s optical properties (e.g., $\delta n/n\approx 1$, where $n$ is the refractive index), together with periodic and coherent ultrafast modulations on timescales comparable to the light cycle, two conditions that remain challenging to achieve simultaneously.
+Consequently, experimental implementations have so far been confined to electrical circuits Reyes-Ayona and Halevi (2015); Wang et al.
+(2023); Lee et al.
+(arXiv:2507.19916), leaving open the challenge of achieving PTCs beyond the microwave regime.
+Here, we demonstrate the first all-optical realization of a PTC, operating at Terahertz (THz) frequencies and achieved with a surface plasmon cavity metamaterial.
+Using a multi-cycle THz-frequency light field of modest amplitude ($\approx 40$kV/cm), we periodically drive the plasmonic metamaterial and achieve strong (near-unity) and coherent (sub-optical-cycle) parametric modulation of its resonance, an effect which arises from dynamical variations in the carriers’ effective mass and reaching up to 80% of their rest mass.
+Via spectroscopic measurements, we uncover the transition to the PTC regime as a function of time which manifests, in particular, as a substantial decrease of the plasmonic losses by more than 50%, opening the door to the temporal engineering of plasmonic losses and plasmonic lasing in this system Kiselev et al.
+(2024b); Feinberg et al.
+(2025); Kiselev et al.
+(2024a).
 Experiment
-To demonstrate our approach, we take a conceptually simple metamaterial structure, namely a periodic lattice of sub-wavelength metal/insulator/plasma cavities, where the plasmonic material corresponds to the small band-gap semiconductor InSb Aupiais_2023 .
+To demonstrate our approach, we take a conceptually simple metamaterial structure, namely a periodic lattice of sub-wavelength metal/insulator/plasma cavities, where the plasmonic material corresponds to the small band-gap semiconductor InSb Aupiais et al.
+(2023).
 It is shown in Fig.
-1 together with the experimental configuration.
+1b together with the experimental configuration.
 Figure 2: Temporal response of the metamaterial under a multicycle drive with a peak field of 40kV/cm: frequency-resolved power reflectivity (central panels) and phase (right panels) spectra as a function of time $\tau$ during the drive (left panels) a) Coarse grained dynamics measured over the complete duration of the drive b) Close up on the dynamics in the vicinity of the peak driving field with sub-cycle temporal resolution.
@@
 The equilibrium power reflectivity of the metamaterial, shown in Fig.
-1b, exhibits an absorption dip centered at $0.77\rm{THz}$ with a full width at half maximum of $\sim 0.18\rm{THz}$.
+1c, exhibits an absorption dip centered at $0.77\rm{THz}$ with a full width at half maximum of $\sim 0.18\rm{THz}$.
 It corresponds to the resonance of the fundamental $\lambda/2$ mode of the sub-wavelength Fabry-Perot surface plasmonic cavity, whose mode profile is displayed in Fig.
-1c.
-The periodically driven regime is investigated using a multi-cycle THz pump pulse centered at $0.69\rm{THz}$—slightly below the equilibrium resonance frequency—with a peak field amplitude of approximately 40kV/cm.
-The system’s dynamics during the drive is monitored using a broadband THz probe pulse, $\tilde{E}(\tau,t)$, where $\tau$ denotes the delay between the pump and the electro-optic sampling beam, and $t$ represents the internal delay between the sampling beam and the THz probe pulse, as defined in Fig.1a.
+1d.
+The periodically driven regime is investigated using a narrowband THz pump pulse centered at $0.69\rm{THz}$—slightly below the equilibrium resonance frequency—with a peak field amplitude of approximately 40kV/cm.
+The system’s dynamics during the drive is monitored by a broadband THz probe pulse, $\tilde{E}(\tau,t)$, where $\tau$ denotes the delay between the pump and the electro-optic sampling beam, and $t$ represents the internal delay between the sampling beam and the THz probe pulse, as defined in Fig.
+1b.
 For each time delay $\tau$, the frequency-resolved response of the metamaterial is obtained by performing a Fourier transform over the internal delay $t$.
@@
 Model
-We attribute the temporal modulations of the surface plasmon cavity resonance to dynamical variations in the carriers’ effective mass $m^{*}(\tau)$ and kinetic inductance associated with non-parabolic conduction bands in small band-gap semiconductors Houver_2019 ; Soranzio_2024 : $m^{*}(\tau)=m^{*}_{0}+\delta m^{*}(\tau)$, where $m^{*}_{0}$ corresponds to the carriers’ rest effective mass.
+We attribute the temporal modulations of the surface plasmon cavity resonance to dynamical variations in its kinetic inductance through the carriers’ effective mass $m^{*}(\tau)$, which we write $m^{*}(\tau)=m_{0}^{*}+\delta m^{*}(\tau)$, where $m_{0}^{*}$ denotes the carriers’ rest effective mass.
+Such variations are known to arise from non-parabolic conduction bands in narrow band-gap semiconductors Houver et al.
+(2019); Soranzio et al.
+(2024) and have also been discussed in the context epsilon-near-zero plasmonic materials operating at optical frequencies Khurgin et al.
+(2020).
 Under periodic driving at frequency $\omega_{\mathrm{d}}$, the electrons’ momentum $p(\tau)$ oscillates as $p(\tau)=p_{0}\cos(\omega_{\mathrm{d}}\tau)$.
-For a non-parabolic and isotropic conduction band given by the Kane model Kane_1957 , this results in effective mass modulations at twice the drive frequency:
-$\delta m^{*}(\tau)=\alpha p^{2}(\tau)=\frac{\alpha}{2}{p_{0}}^{2}+\frac{\alpha}{2}{p_{0}}^{2}\cos(2\omega_{\mathrm{d}}t)$ (1)
+For a non-parabolic and isotropic conduction band given by the Kane model Kane (1957), this results in effective mass modulations at twice the drive frequency:
+$\delta m^{*}(\tau)=\alpha p^{2}(\tau)=\frac{\alpha}{2}{p_{0}}^{2}+\frac{\alpha}{2}{p_{0}}^{2}\cos(2\omega_{\mathrm{d}}\tau)$ (1)
 where $\alpha$ accounts for the deviation from a purely parabolic band dispersion ($\alpha>0$ here, see methods).
@@
 Using an LC-resonator model for the surface plasmonic cavity, the resonance frequency is given by $\omega_{\mathrm{c}}=\frac{1}{\sqrt{LC}},$ where $L$ and $C$ are the inductance and capacitance of the mode, respectively.
-For surface plasmons, the inductance $L=L_{G}+L_{K}$ includes contributions from both a geometric inductance $L_{G}$, arising from retardation effects, and a kinetic inductance $L_{K}\propto\frac{m^{*}}{ne^{2}}$, which represents part of the mode energy that is stored into the kinetic energy of the carriers (with $n$ the carrier density and $e$ the charge) Staffaroni_2012 .
+For surface plasmons, the inductance $L=L_{G}+L_{K}$ includes contributions from both a geometric inductance $L_{G}$, arising from retardation effects, and a kinetic inductance $L_{K}\propto\frac{m^{*}}{ne^{2}}$, which represents part of the mode energy that is stored into the kinetic energy of the carriers (with $n$ the carrier density and $e$ the charge) Staffaroni et al.
+(2012).
 The kinetic inductance can be separated into an equilibrium and a time-varying part: $L_{K}=L_{K}^{0}+\delta L_{K}(\tau)$.
@@
 where the time-dependent modulation parameter is given by $\eta(\tau)=-\frac{\omega_{\rm c}^{0}}{2}\frac{L_{K}^{0}}{L_{G}+L_{K}^{0}}\beta\frac{\delta m^{*}(\tau)}{m^{*}_{0}}$ (see methods and supplementary material).
-This Hamiltonian describes a parametrically driven cavity resonance and the emergence of an interaction between surface plasmons that is generated by the drive.
 To capture the metamaterial response function that is measured experimentally -the reflectivity-, we implement an input-output formalism with two baths representing radiative and non-radiative (ohmic) decay channels.
@@
 Adopting the experimental temporal conventions and using a single adjustable parameter, we present in Fig.
-2c the computed time-resolved reflection coefficient $r(\tau,\nu)$ that best matches the experimental observations.
+2c the computed time-resolved reflection coefficient $\tilde{r}(\tau,\omega)$ that best matches the experimental observations.
 This approach allows direct comparison between experimental and theoretical response functions, enabling inference of the time-dependent modulation of microscopic system parameters.
+Figure 3: a) Real and imaginary parts of the zero-order (center panel) and first-orders (outer panels) Floquet spectra obtained from Fourier Transform of the time domain data of Fig 2b.
+over one drive period at $\tau\approx 24ps$.
+Experimental data points are shown together with a fitting of the lineshapes (see text).
+b) Evolution of the complex-frequency poles of the reflectivity, controlled by the Floquet eigenmodes, as a function of drive strength across the PTC transition.
+c) Real and imaginary parts of $\Delta\tilde{\mu}/\omega_{c}^{0}$ as a function of drive strength $\eta_{0}/\omega_{c}$ computed for parameters taken from the experiment: $\omega_{d}/\omega_{c}^{0}=0.92$, $\Gamma/\omega_{c}^{0}=0.12$.
+Dashed lines indicate the values of $\eta_{0}/\omega_{c}$ taken for the computation of the spectra shown in Fig.
+3d.
+d) Spectra of the real part of $\tilde{r}_{0}(\omega)$ and imaginary part of $\tilde{r}_{-1}(\omega)$ across the PTC transition and computed for values of $\eta_{0}/\omega_{c}$ indicated by the dashed lines in panel c.
 Discussion
@@
 This corresponds to an impressive $\delta m^{*}(\tau)/m^{*}_{0}\approx 80\%$ dynamical effective mass modulations.
-As a function of driving strength, we further establish that these parametric effects are saturated for the reported peak driving field of 40kV/cm.
-This is indicated by the leveling-off of the response and the emergence of incoherent nonlinearities measured at a larger field of $140\textrm{kV/cm}$ (see supplementary material).
-Hence, this value appears to be a physical bound for coherent effective mass modulations in this system.
-Given this value, a strategy to enhance further parametric effects consists in increasing the ratio $L_{K}^{0}/L_{G}$.
-It can be achieved by decreasing the size of the plasmonic cavities so as to operate in the ’electrostatic’ regime where retardation effects vanish and kinetic ones dominate Staffaroni_2012 ; Greffet_2012 : $L_{K}^{0}\gg L_{G}$ and $\beta=1$.
-In the following, using Floquet analysis with periodic drive $\eta(\tau)=-\eta_{0}\cos^{2}(\omega_{\mathrm{d}}\tau)$ and accounting for the observed losses, we highlight three key physical phenomena lying within the experimentally accessible range of parameters of this system for : 1) the present cavity design ($\beta\approx 0.65,L_{K}^{0}/L_{G}\approx 2.4$, $\eta_{0}/\omega_{\mathrm{c}}^{0}\approx 0.18$) and 2) an electrostatic plasmonic cavity ($\beta=1,L_{K}^{0}\gg L_{G}$, $\eta_{0}/\omega_{\mathrm{c}}^{0}=0.4$).
-Parametric amplification of surface plasmons
-As well known from other parametrically driven systems, the modulation induced by the drive generates gain on the cavity mode.
-The corresponding growth rate of the mode $\mathrm{Im}(\mu)/\omega_{\mathrm{c}}^{0}$, with $\mu$ being an eigenvalue of the Floquet matrix describing long-term behavior of the system (see methods), is reported in the phase diagram of Fig.
-3a as a function of drive frequency and modulation strength, together with the operating points of the two aforementioned cavity designs.
-When the growth rate overcomes the total loss rate $\Gamma/\omega_{\mathrm{c}}^{0}$ (including radiative and non-radiative contributions), the system develops a parametric instability: it crosses a threshold illustrated by the solid contour line and defined by $\mathrm{Im}(\mu)/\omega_{\mathrm{c}}^{0}=\Gamma/\omega_{\mathrm{c}}^{0}$ in Fig.
-3a for which quantities diverge exponentially and parametric amplification sets in.
-Even before reaching this threshold and within the broader region defined by the dashed contour line, the system remains stable as a function of time but demonstrates average reflectivities above 1 (see methods and supplementary material).
-This highlights the strong potential of this platform for achieving light amplification and lasing of surface plasmons.
-Figure 3: Plasmonic time crystal phenomena computed for the experimentally accessible regime of parameters of this platform
-a) Stability diagram: growth rate of the cavity mode as a function of drive frequency $\omega_{\mathrm{d}}$ and modulation strength $\eta_{0}$ relative to the equilibrium cavity frequency $\omega_{\mathrm{c}}^{0}$.
-The solid contour line defined by $\mathrm{Im}(\mu)/\omega_{\mathrm{c}}^{0}=\Gamma/\omega_{\mathrm{c}}^{0}=0.12$ marks the threshold for parametric amplification given the observed losses.
-Dashed contour line indicate regions exhibiting average reflectivity above one.
-Operating points of the present cavity design (star) and an electrostatic cavity (dot) are indicated (see text) b) Wigner function of the cavity state accounting for the losses.
-Top panel: no drive.
-Middle panel: this experiment’s parameters ($\eta_{0}/\omega_{\mathrm{c}}^{0}=0.18$, $\omega_{\rm d}/\omega_{\mathrm{c}}^{0}=0.92$) Bottom panel: an electrostatic cavity ($\eta_{0}/\omega_{\mathrm{c}}^{0}=0.4$, $\omega_{\rm d}/\omega_{\mathrm{c}}^{0}=0.8$) c Plasmonic time crystal $k$-gap openings.
-Real and imaginary parts of the eigenmodes’ frequency $\omega$ as a function of momentum (or Bloch-momentum) $k$, normalized to the plasma frequency $\omega_{\mathrm{p}}$ and including the losses, for the equilibrium (dashed lines) and the temporally-modulated systems (solid lines).
-Top panel: a dispersive insulator/plasmonic interface.
-Bottom panel: our dispersion-free plasmonic metamaterial design for cavities tuned in the electrostatic regime (shown in the first Brillouin zone of the metamaterial of period $d$.
-The drive frequency corresponds to $\omega_{\mathrm{d}}/\omega_{\mathrm{p}}=0.68$ in both cases.
-Generation of squeezed and entangled plasmons
-[…]
-$k$-agnostic gap openings in plasmonic metamaterial time crystals
-Moving from a single cavity picture to the actual periodically structured metamaterial, a central aspect of time crystals lies in the opening of $k$-gaps in their dispersion relations Asgari_2024 , where $k$ denotes the momentum (translational invariant systems) or quasi-momentum (periodically structured metamaterials) of photonic eigenmodes.
-Since material dispersion limits the width of the $k$-gaps, several strategies were recently proposed to widen their span Wang_2024 ; Feinberg_2025 .
-The ability of our metamaterial approach to engineer a flat dispersion allows to address such challenges Aupiais_2023 .
-In Fig 3c.
-we compare a dispersive insulator/plasmonic interface together with our metamaterial design and compute $k$-gap openings in the dispersion relation (see methods).
-For the dispersive plasmonic interface (top panel), although the $k$-gaps are already significant, they fail to cover the whole dispersion.
-In contrast, our dispersion-free metamaterial (bottom panel) achieves $k$-gap openings across the full Brillouin zone, effectively realizing a $k$-agnostic time crystal.
+This value is not limited by the maximum driving field that can be applied, which could be further increased, but rather by the onset of incoherent nonlinearities, namely impact ionization (see Supplementary Material), which starts to emerge at fields exceeding $\sim 40kV/cm$ due to the small band-gap of the InSb plasmonic material Houver et al.
+(2019); Biasco et al.
+(2022).
+Hence, it appears to be a physical bound for coherent effective mass modulations in this system.
+Spectroscopy of a PTC
+We next investigate the periodically driven quasi–steady-state regime observed at $\tau_{0}\simeq 24\,\mathrm{ps}$ by analyzing the dynamics over one period of the drive $\tau\in[\tau_{0},\tau_{0}+2\pi/\omega_{d}]$ within the framework of Floquet theory.
+For a system that is periodic in time with frequency $2\omega_{d}$, the complex reflectivity $\tilde{r}(\tau,\omega)$ can be expanded in Fourier harmonics at frequencies $\pm 2n\omega_{d}$ (see Supplementary material).
+In the present case, this expansion is dominated by the zeroth-order (time-averaged) component and the first-order harmonic contributions, such that:
+$\tilde{r}(\tau,\omega)=\tilde{r}_{b}(\omega)+\tilde{r}_{0}(\omega)+\sum_{n=\pm 1}\tilde{r}_{n}(\omega)\,e^{2in\omega_{d}\tau}.$ (4)
+The time-averaged response consists of a non-resonant background $\tilde{r}_{b}(\omega)$ and a resonant cavity contribution $\tilde{r}_{0}(\omega)$, while $\tilde{r}_{\pm 1}(\omega)$ are the first-order Floquet sidebands.
+The latter correspond to photons whose frequency has been either up-converted, $\tilde{r}_{1}(\omega)$, or down-converted, $\tilde{r}_{-1}(\omega)$, by $2\omega_{d}$ with respect to their incident frequency $\omega$.
+Spectra for $\tilde{r}_{0}(\omega)$ and $\tilde{r}_{\pm 1}(\omega)$ are presented in Fig.
+3a.
+Although the reflectivity satisfies $\tilde{r}(\tau,\omega)=\tilde{r}(\tau,-\omega)^{*}$ at any time $\tau$, spectra of the Floquet sidebands $\tilde{r}_{\pm 1}(\omega)$ display two striking features: as opposed to $\tilde{r}_{0}(\omega)$, they are predominantly one-sided in frequency and their lineshapes deviate qualitatively from that expected of a simple resonance.
+Instead, they manifest characteristics of a second order resonance (or pole) emanating from the presence of two closely spaced resonances within the spectral linewidth (see below).
+As shown in Fig.
+3b, these aspects can be interpreted as follows.
+In the absence of modulation, the cavity resonates for both positive ($\omega\approx\omega_{c}^{0}$) and negative ($\omega\approx-\omega_{c}^{0}$) frequencies within the spectral width $\Gamma$, which is described mathematically as two poles in the reflectivity located in the complex-frequency plane at $\tilde{\omega}_{c}^{0}=\omega_{c}^{0}-i\Gamma$ and $(\tilde{\omega}_{c}^{0})^{*}=-\omega_{c}^{0}-i\Gamma$ (Fig.
+3b-panel (1)).
+The periodic drive generates Floquet replicas of these poles shifted by integer multiples of $\pm 2\omega_{d}$.
+Notably, when the drive frequency is tuned close to the cavity resonance ($\omega_{d}\approx\omega_{c}^{0}$), the first-order replica of the negative (resp.
+positive) frequency component becomes nearly degenerate with its original positive (resp.
+negative) frequency counterpart i.e., $\omega_{c}^{0}+2\omega_{d}\approx\omega_{c}^{0}$ (resp.
+$\omega_{c}^{0}-2\omega_{d}\approx-\omega_{c}^{0}$) (Fig.
+3b-panel (2)).
+These two Floquet eigenmodes are described as two closely spaced poles, which can be written on the positive frequency side as:
+$\tilde{\mu}_{\pm}=\tilde{\omega}_{d}\pm\Delta\tilde{\mu}$, (5)
+where $\tilde{\omega}_{d}=\omega_{d}-i\Gamma$.
+One can show that the Floquet matrix displays PT symmetry (see Supplementary material) which, if unbroken, constrains $\Delta\tilde{\mu}$ to be real-valued Bender and Boettcher (1998); El-Ganainy et al.
+(2018), as it is the case for weak driving, see Fig.
+3c.
+In this regime, photons undergo repeated scattering between the positive and negative frequency components of the resonance and find a resonant pathway to get either up or down converted.
+Importantly, non-resonant pathways in which, for instance, positive frequency photons attempt to get up-converted are strongly inhibited, explaining why the Floquet sidebands appear predominantly one-sided (e.g.
+$|\tilde{r}_{1}(+\omega)|\ll|\tilde{r}_{1}(-\omega)|$).
+Ultimately, the Floquet sidebands probe the joint spectral density of states, i.e.
+the spectral overlap $\propto\frac{1}{(\omega-\tilde{\mu}_{+})(\omega-\tilde{\mu}_{-})}$, of the two Floquet eigenmodes, which naturally accounts for their second-order pole character (see Supplementary material).
+As the driving strength is increased, the coupling between those two eigenmodes gets stronger, their separation $\Delta\tilde{\mu}$ decreases and eventually vanishes, at which point they coalesce at $\tilde{\omega}_{d}$ (Fig.
+3b-panel (3)).
+This coalescence defines an exceptional point characterized by a spontaneous breaking of PT symmetry Koutserimpas et al.
+(2018); Bossart and Fleury (2021) and marks the transition into the PTC regime.
+Beyond this point, the difference between the two Floquet eigenmodes $\Delta\tilde{\mu}$ become purely imaginary
+$\tilde{\mu}_{\pm}=\tilde{\omega}_{d}\pm i\Delta\Gamma$ (6)
+leading to reduced damping in one of the two modes, which is a hallmark of the PTC phase (Fig. 3b-panel (4)) Asgari et al. (2024); Feinberg et al. (2025).
+In Fig.
+3c-d, using Floquet analysis with periodic drive $\eta(\tau)=-\eta_{0}\cos^{2}(\omega_{\mathrm{d}}\tau)$, we compute the evolution of $\Delta\tilde{\mu}/\omega_{c}^{0}$ (Fig.
+3c) and of the spectral lineshapes expected across this transition (Fig 3d), from which two important conclusions can be drawn: 1) the spectra evolve smoothly across the PTC transition and 2) the lineshapes’ characteristics of $\tilde{r}_{0}(\omega)$ and $\tilde{r}_{\pm 1}(\omega)$ (both real and imaginary parts) encode detailed information about the poles’ structure and the system’s parameters, notably the spectral locations of $\tilde{\mu}_{\pm}$ which can be obtained through fitting of the experimental lineshapes (see Methods and Supplementary material).
+Figure 4: a Left panel: Temporal evolution of the Floquet eigenvalues $\tilde{\mu}_{\pm}$ (solid lines) shown with 95% confidence interval Right panel: Evolution of the Floquet eigenvalues $\tilde{\mu}_{\pm}$ as a function of drive strength $\eta_{0}/\omega_{c}^{0}$ computed from the model for parameters taken from the experiment.
+c) Stability diagram of the driven system: parametric gain $\Delta\Gamma$ as a function of drive frequency $\omega_{\mathrm{d}}$ and modulation strength $\eta_{0}$ relative to the equilibrium cavity frequency $\omega_{\mathrm{c}}^{0}$.
+The colored area where $\Delta\Gamma\neq 0$ signals the PTC regime.
+The solid contour line defined by $\Gamma/\omega_{\mathrm{c}}^{0}=0.12$ marks the threshold for plasmonic lasing given the measured equilibrium losses $\Gamma$.
+Temporal transition to the PTC regime
+Below, we unambiguously confirm the transition to the PTC regime by extracting the spectral locations of the Floquet eigenmodes as a function of time $\tau$.
+Treating the drive amplitude as a slowly varying function of time with respect to the fast modulations at $\pm 2\omega_{d}$, this is achieved by performing a short term Fourier Transform of the time-domain data (Fig.
+2b) to extract the slowly varying components $\tilde{r}_{0}(\tau,\omega)$ and $\tilde{r}_{1}(\tau,\omega)$ and performing a fit of the lineshapes to extract the spectral location of the poles.
+The results are presented in Fig 4a.
+As a function of time, we observe the pole coalescence and the transition to the PTC regime at around $\tau\approx 21ps$ before the poles acquire different imaginary part.
+The experimental results are seen to be in excellent agreement with predictions of our model, notably in the magnitude of the effect obtained.
+Within the PTC regime, the emergence of gain is directly reflected in the reduction of the imaginary part of one of the poles.
+Experimentally, this corresponds to a linewidth narrowing of approximately $\Delta\Gamma\approx 40GHz$ from an initial equilibrium linewidth $\Gamma=\gamma_{r}+\gamma_{nr}$, with $\gamma_{nr}\approx 70GHz$ and $\gamma_{r}\approx 20GHz$ being the non radiative and radiative losses respectively.
+It indicates that the non-radiative plasmonic losses $\gamma_{nr}$ are reduced by more than 50%.
+To the best of our knowledge, this constitutes the first demonstration of plasmonic loss reduction induced by a parametric drive, opening a new route for plasmonic loss engineering based on a temporal approach Kiselev et al.
+(2024b); Feinberg et al.
+(2025).
+Towards plasmonic lasing
+Finally, we place these results in the broader context of this platform’s parameters and discuss the possibility of plasmonic lasing in this system, which may occur when the gain produced by the drive overcomes the total losses ($\Delta\Gamma>\Gamma$).
+To that end, we compute in Fig.
+4b its phase diagram as a function of drive frequency and strength.
+The colored area marks the PTC phase with the emergence of gain ($\Delta\Gamma\neq 0$), while the white area corresponds to the absence of gain ($\Delta\Gamma=0$).
+The current experimental realization of Fig 4a are reported as colored dots on this diagram.
+We predict that plasmonic lasing with this platform is within experimental reach with an improved cavity design and fine tuning of the drive parameters.
+A strategy for this consists in increasing the ratio $L_{K}^{0}/L_{G}$ by decreasing the size of the plasmonic cavities so as to operate in the ’electrostatic’ regime where retardation effects vanish and kinetic ones dominate Staffaroni et al.
+(2012); Greffet (2012).
+For such an electrostatic cavity ($\beta=1,L_{K}^{0}\gg L_{G}$, $\eta_{0}/\omega_{\mathrm{c}}^{0}=0.4$, white dot in Fig 4b), this enhances parametric effects as compared to the present cavity design ($\beta\approx 0.65,L_{K}^{0}/L_{G}\approx 2.4$, $\eta_{0}/\omega_{\mathrm{c}}^{0}\approx 0.18$, star in Fig 4b), and may ultimately allow to cross the threshold for lasing, as indicated by the white solid line shown in Fig.
+4.
 Conclusion
-In conclusion, we report the first optical implementation of a photonic time crystal.
-This is realized in a plasmonic system and enabled by the sizeable and coherent dynamical variations of the carriers’ effective mass.
-Our work establishes that this mechanism, previously known in bulk systems, can be extended to surface plasmons and leveraged as a foundational element for constructing plasmonic metamaterials harbouring Floquet physics and time crystalline phenomena.
-Along those lines, the inherent tunability of the plasmonic medium to external parameters Aupiais_2023 ; Aupiais_2024 offers a unique route towards extension of PTC phenomena to other photonic degrees of freedom, including spin and orbital ones.
+In conclusion, we report the first all-optical realization of a photonic time crystal, operating at THz frequencies and implemented with a plasmonic metamaterial.
+The underlying mechanism relies on large and ultrafast dynamical modulations of the carriers’ effective mass sustaining the plasmonic resonance.
+The transition to the PTC regime, evidenced by direct spectroscopic measurements and fully supported by a theoretical analysis based on Floquet theory, manifests through the coalescence of two Floquet-driven optical modes followed by the onset of plasmonic gain.
+This work paves the way towards the temporal engineering of dissipation and light-matter interactions in plasmonic systems and establishes a solid framework for the realization and spectroscopic identification of PTC phenomena in future implementations.
 Acknowledgments
@@
 The top metallic pattern is then defined through photolithography, followed by Ti/Au (20 nm/200 nm) evaporation and lift-off.
-The resulting metamaterial structure consists of periodic metallic stripes with a width of $s=41\penalty 10000\ \mu\mathrm{m}$ and spacing $a=16\penalty 10000\ \mu\mathrm{m}$, yielding a total period of $d=s+a=57\penalty 10000\ \mu\mathrm{m}$ and follows the design reported in a previous work Aupiais_2023 .
+The resulting metamaterial structure consists of periodic metallic stripes with a width of $s=41\penalty 10000\ \mu\mathrm{m}$ and spacing $a=16\penalty 10000\ \mu\mathrm{m}$, yielding a total period of $d=s+a=57\penalty 10000\ \mu\mathrm{m}$ and follows the design reported in a previous work Aupiais et al.
+(2023).
 THz pump-probe spectroscopy
@@
 $C$ and $L$ parametrize the capacitive and inductive parts of the energy of the plasmon field.
-The inductance $L$ has two contributions Staffaroni_2012 :
+The inductance $L$ has two contributions Staffaroni et al. (2012):
   • a geometric inductance $L_{G}$ arising from retardation effects of the electromagnetic field
@@
 Computation of the parameters entering equations (11) and (12) is outlined here and explained in more detail in the supplementary material.
-These calculations are based on the determination of the cavity resonance frequency $\omega_{\rm c}$, obtained by solving the dispersion relation of a truncated metal/insulator/plasmonic waveguide Aupiais_2023 , when parameters are varied.
+These calculations are based on the determination of the cavity resonance frequency $\omega_{\rm c}$, obtained by solving the dispersion relation of a truncated metal/insulator/plasmonic waveguide Aupiais et al.
+(2023), when parameters are varied.
 Notably, the resonance frequency depends, among other geometrical and optical parameters of the constitutive elements of the metamaterial, on the plasma frequency of the plasmonic medium supporting the resonance: $\omega_{\rm p}=\sqrt{\frac{ne^{2}}{\epsilon_{0}\epsilon_{\infty}m^{*}}}$, where $n$ and $e$ are the density and charge of the carriers, $m^{*}$ their effective mass, $\epsilon_{0}$ the vacuum permittivity and $\epsilon_{\infty}$ the high-frequency permittivity of the plasmonic medium.
@@
 Similarly, the dependence of kinetic inductance with the effective mass, allowing to determine $\delta L_{K}$ and $\beta$, is obtained through a variation of the effective mass that enters the plasma frequency, all other parameters being fixed.
-We point out that these computations are parameter-free and obtained from experimental measurements of the geometrical and optical properties of the constitutive elements of the metamaterial, as detailed in Aupiais_2023 .
+We point out that these computations are parameter-free and obtained from experimental measurements of the geometrical and optical properties of the constitutive elements of the metamaterial, as detailed in Aupiais et al.
+(2023).
 Temporal dynamics (Figure 2)
@@
 2, we determine $\eta(\tau)$ from equation (12) as follows.
-We consider a non-parabolic band structure given by the Kane model suitable for the description of small-band gap semiconductors Kane_1957 where the electronic dispersion relation $\varepsilon(p)$ satisfies $\varepsilon(p)(1+\alpha\varepsilon(p))=\frac{p^{2}}{2m^{*}_{0}}$ and $\alpha$ is the parameter that accounts for non-parabolicity.
-The effective mass $m^{*}(p)$ is defined by $m^{*}(p)=p\left(\frac{d\varepsilon}{dp}\right)^{-1}$, which allows to express it as: $m^{*}(p)=m^{*}_{0}(1+2\alpha\varepsilon(p))$ Zawadzki_1974 .
+We consider a non-parabolic band structure given by the Kane model suitable for the description of small-band gap semiconductors Kane (1957) where the electronic dispersion relation $\varepsilon(p)$ satisfies $\varepsilon(p)(1+\alpha\varepsilon(p))=\frac{p^{2}}{2m^{*}_{0}}$ and $\alpha$ is the parameter that accounts for non-parabolicity.
+The effective mass $m^{*}(p)$ is defined by $m^{*}(p)=p\left(\frac{d\varepsilon}{dp}\right)^{-1}$, which allows to express it as: $m^{*}(p)=m^{*}_{0}(1+2\alpha\varepsilon(p))$ Zawadzki (1974).
 Keeping only the leading order term $\propto p^{2}$ in the electronic dispersion relation, we parametrize the effective mass as $m^{*}(p)=m^{*}_{0}(1+2\alpha\frac{p^{2}}{2m^{*}_{0}})$, so that $\frac{\delta m^{*}(p)}{m^{*}_{0}}=\frac{\alpha p^{2}}{m^{*}_{0}}$.
@@
 Parameter Name Notation Value (SI Unit) Source
-Non-parabolicity parameter $\alpha$ $5.6\penalty 10000\ \mathrm{eV}^{-1}$ Zawadzki_1974
+Non-parabolicity parameter $\alpha$ $5.6\penalty 10000\ \mathrm{eV}^{-1}$ Zawadzki (1974)
 Radiative power decay rate $\gamma$ $2\pi\times 0.042\penalty 10000\ \mathrm{THz}$ Experimental fit
@@
 Inductance to effective mass variations $\beta$ $0.65$ Computed
-Floquet analysis (Figure 3)
-For the Floquet analysis presented in Fig. 3, we use a periodic drive: $\eta(\tau)=-\eta_{0}\cos^{2}(\omega_{\mathrm{d}}\tau)$.
-Phase diagram (Figure 3a)
-Under a periodic drive $\eta(t)$, the evolution operator of the cavity mode (a^,a^)($\hat$$^{\dagger},\hat{a}$) is decomposed using Floquet’s theorem as:
+Spectroscopic analysis based on Floquet theory (Figures 3 & 4)
+Floquet theory
+Under a drive $\eta(\tau)=-\eta_{0}\cos^{2}(\omega_{d}\tau)$ which is periodic with frequency $2\omega_{d}$, the evolution operator of the cavity mode (a^,a^)($\hat$$^{\dagger},\hat{a}$) is decomposed using Floquet’s theorem as:
 $\mathbf{U}(\tau)=\mathbf{G}(\tau)e^{-\mathrm{i}\mathbf{\mu}\tau},$ (14)
-where $\mathbf{G}(\tau)$ and $\mathbf{\mu}$ are both $2\times 2$ matrices acting on the spinor space of the cavity mode (a^,a^)($\hat$$^{\dagger},\hat{a}$) (see supplementary material).
-$\mathbf{G}(\tau)$ is periodic with the same period as the drive and describes the periodic micro-motion of the system while the Floquet exponent matrix $e^{-\mathrm{i}\mathbf{\mu}t}$ describes its long-term behavior.
-In the absence of dissipation, the imaginary part of the eigenvalues $\mu_{1,2}$ of matrix $\mathbf{\mu}$ determine wether the system is stable ($\mathrm{Im}(\mu_{1,2})<0$) or unstable ($\mathrm{Im}(\mu_{1,2})>0$), as quantities evolve as $e^{\mathrm{Im}(\mu_{1,2})t}$ at long times.
-In the presence of dissipation characterized by a total decay rate $\Gamma=\frac{\gamma+\gamma^{\prime}}{2}$, it can be shown that eigenvalues of $\mathbf{\mu}$ undergo a shift by $\Gamma$ while $\mathbf{G}(t)$ remains unaffected.
-Defining $\mathrm{Im}(\mu)=\max\limits_{1,2}\mathrm{Im}(\mu_{1,2})$, this allows to conclude directly on the stability of the dynamics from a contour line defined by $\frac{|\mathrm{Im}(\mu)|}{\omega_{\mathrm{c}}^{0}}=\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$ and which marks the transition between stable ($\frac{|\mathrm{Im}(\mu)|}{\omega_{\mathrm{c}}^{0}}<\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$) and unstable dynamics ($\frac{|\mathrm{Im}(\mu)|}{\omega_{\mathrm{c}}^{0}}>\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$).
+where $\mathbf{G}(\tau)$ and $\mathbf{\mu}$ are both $2\times 2$ matrices acting on the spinor space of the cavity mode (a^,a^)($\hat$$^{\dagger},\hat{a}$).
+$\mathbf{G}(\tau)$ is periodic with the same period as the drive and describes the periodic micro-motion of the system while the Floquet exponent matrix $e^{-\mathrm{i}\mathbf{\mu}\tau}$ describes its long-term behavior.
+In the supplementary material, it is shown that the reflection coefficient can be decomposed in Fourier series as:
+$\tilde{r}(\tau,\omega)=\tilde{r}_{b}(\omega)+\sum_{n}\tilde{r}_{n}(\omega)e^{\mathrm{i}2n\omega_{d}\tau}$ (15)
+where $\tilde{r}_{b}(\omega)$ corresponds to a non-resonant background contribution to the reflectivity and the $n^{th}$ order Fourier harmonic $\tilde{r}_{n}(\omega)$ can be expressed as:
+$\frac{\tilde{r}_{n}(\omega)}{-\mathrm{i}\gamma}=\penalty 10000\ \frac{\tilde{c}_{n}(\omega-\tilde{\omega}_{d})+\tilde{d}_{n}}{(\omega-\tilde{\mu}_{+})(\omega-\tilde{\mu}_{-})}$ (16)
+Here, $\gamma$ corresponds to the radiative power decay rate, the Floquet eigenvalues $\tilde{\mu}_{\pm}$ are defined through $\tilde{\mu}_{\pm}=\tilde{\omega}_{d}\pm\Delta\tilde{\mu}$ with $\tilde{\omega}_{d}\overset{\text{def}}{=}\omega_{d}-i\Gamma$ and can be obtained through diagonalization of the Floquet matrix $\mathbf{\mu}$, while the coefficients $(\tilde{c}_{n},\tilde{d}_{n})$ are related to the Fourier coefficients of the $\mathbf{G}(\tau)$ matrix.
+Their complete expression can be found in the supplementary material.
+Spectral lineshapes (Figure 3a,d)
+The experimental lineshapes of $\tilde{r}_{0}(\omega)$ and $\tilde{r}_{\pm 1}(\omega)$ are obtained from a short term Fourier transform of the time domain data (Fig.
+2b) performed over a sliding window $\tau,\tau+2\pi/\omega_{d}]$ as a function of time $\tau$.
+They are presented in Figure 3a for $\tau\approx 24ps$.
+The complete temporal evolution of $\tilde{r}_{0}(\omega)$ and $\tilde{r}_{\pm 1}(\omega)$ is reported in the Supplementary material.
+Theoretically, they are obtained from numerical computation of formula (16).
+Experimental determination of the Floquet eigenvalues (Figure 4a)
+The Floquet eigenvalues presented in Fig.
+4a are obtained from fitting the experimental lineshapes according to equation (16).
+Other coefficients are reported in the supplementary material.
+The fit is simultaneously performed on the real and imaginary parts of $\tilde{r}_{0}(\omega)$ and $\tilde{r}_{-1}(\omega)$ together.
+As shown in the supplementary material, the fitting procedure is benchmarked on theoretical data to ensure that it accurately reproduces the eigenvalues that can obtained independently from diagonalization of the Floquet matrix.
+Stability diagram (Figure 4b)
+In the absence of dissipation, the imaginary part of the eigenvalues $\mu_{\pm}$ determine wether the system is stable ($\mathrm{Im}(\mu_{\pm})<0$) or unstable ($\mathrm{Im}(\mu_{\pm})>0$), as quantities evolve as $e^{\mathrm{Im}(\mu_{\pm})\tau}$ at long times.
+In the presence of dissipation characterized by a total decay rate $\Gamma=\frac{\gamma+\gamma^{\prime}}{2}$, it can be shown that eigenvalues of $\mathbf{\mu}$ undergo a shift by $i\Gamma$ while $\mathbf{G}(t)$ remains unaffected.
+This allows to conclude directly on the stability of the dynamics from a contour line defined by $\frac{|\Delta\Gamma|}{\omega_{\mathrm{c}}^{0}}=\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$ and which marks the transition between stable ($\frac{|\Delta\Gamma|}{\omega_{\mathrm{c}}^{0}}<\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$) and unstable dynamics ($\frac{|\Delta\Gamma|}{\omega_{\mathrm{c}}^{0}}>\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$).
 The total normalized decay rate $\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}=0.12$ presented in Fig.
-3a.
+4b.
 corresponds to the experimentally measured one and can be found in Table 1.
-Wigner function (Figure 3b)
-The Wigner function distribution of the cavity state, plotted in Fig. 3b., is defined via its density matrix $\hat{\rho}(\tau)$ as :
-$W(\alpha,\tau)=\frac{2}{\pi}\text{Tr}\left(\hat{\rho}(\tau)\hat{D}(\alpha)(-1)^{a^{\dagger}a}\hat{D}(-\alpha)\right),$ (12)
-where for all $\alpha\in\mathbb{C}$, the displacement operator is defined by $\hat{D}(\alpha)=\exp(\alpha a^{\dagger}-\alpha^{\ast}a)$.
-It is computed from the time evolution of the density matrix according to its Hamiltonian eq.(7) and accounting for losses at the rate $\Gamma=\frac{\gamma+\gamma^{\prime}}{2}$ as :
-$\partial_{t}\hat{\rho}=-\mathrm{i}\left[\hat{H}(\tau),\hat{\rho}\right]+\frac{\gamma+\gamma^{\prime}}{2}\mathcal{D}[a]\hat{\rho}$. (13)
-where $\mathcal{D}[a]$ is the the one-photon loss dissipator which acts on the density matrix through $\mathcal{D}[a]\rho=a\rho a^{\dagger}-\frac{1}{2}\left\{a^{\dagger}a,\rho\right\}$.
-Further details of the calculations are given in the supplementary materials.
-Photonic time crystal (Figure 3c)
-To treat the case of photonic time crystals, we extend the formulation of the model to general dispersive systems.
-Eigenmodes depend on index $k$, which corresponds either to the parallel wavevector (translationally-invariant systems) or Bloch-wavevector (periodically structured metamaterials) of a given eigenmode.
-For such systems, the Hamiltonian writes (see supplementary materials):
-$\hat{H(\tau)}=\sum_{k}\hbar(\omega_{\rm c}^{0}(k)+\eta_{k}(\tau))\hat{a}_{k}^{\dagger}\hat{a}_{k}-\frac{\hbar\eta_{k}(\tau)}{2}\left(\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}+{\hat{a}}_{k}{\hat{a}}_{-k}\right)$ (14)
-where $\omega_{\rm c}^{0}(k)$ and $\eta_{k}(\tau)$ are, respectively, the $k$-dependent equilibrium resonance frequency and modulation parameter of the eigenmodes.
-As we consider systems that are driven homogeneously at $k=0$, creation and annihilation of plasmon pairs have opposite momenta in order to satisfy momentum conservation rules.
-Computation of $\omega_{\rm c}^{0}(k)$ and $\eta_{k}(\tau)$ is performed with an identical procedure as described above for a single cavity: the dispersion relation of the system is computed both at equilibrium and through variation of the effective mass that enters the plasma frequency, allowing for the determination of the relation of proportionality as in eq.
-9 between $\frac{\eta_{k}(\tau)}{\omega_{\rm c}^{0}(k)}$ and $\frac{\delta m^{*}(\tau)}{m_{0}^{*}}$ for any given $k$-mode.
-Floquet analysis is then performed on Hamiltonian eq.
-14 in order to determine the Floquet eigenvalues $\omega_{1,2}$ of the driven system and to track the opening of $k$-gaps.
-After accounting for non-radiative losses with a shift in the imaginary part and normalizing by the equilibrium plasma frequency $\omega_{\mathrm{p}}$, real and imaginary parts are reported in Fig.
-3c in the two following cases for: 1) a dispersive insulator/plasmonic interface and 2) our dispersion-free metamaterial design for cavities tuned in the electrostatic regime.